There are a total of 108 foreign language students in a high school where they offer Spanish, French, and German.
There are 21 students who take at least 2 languages at once.
If there are
44 students of Spanish,
45 students of French, and
40 students of German,
how many students take all three languages at once?

Answer by edjones(8007)

Let 21 take spanish and french.
Then 44-21=23 take only spanish.
45-21=24 take only french.
40 take only German.
21+23+24+40=108
None take all 3 languages.
.
Ed